OpenStudy (anonymous):

Simplify. Express the quotient as a rational exponent. x to the 4 fifths power over x to the 1 third power

OpenStudy (ash2326):

\[\Huge \frac{x^{\frac{4}{5}}}{x^{\frac 1 3}}\] we know that \[\large \frac{a^b}{a^c}=a^{b-c}\] can you apply this here @evergirl?

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

ahhhhhhhhhh................i dunno

OpenStudy (ash2326):

try it subtract 1/3 from 4/5 and put it as power of x

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

were you able to simplify?